### v1

# Survey on Large Language Model-Enhanced Reinforcement Learning: Concept, Taxonomy, and Methods

Abstract
With extensive pre-trained knowledge and high-level general capabilities, large language models (LLMs) emerge as a promising avenue to augment reinforcement learning (RL) in aspects such as multi-task learning, sample efficiency, and task planning. In this survey, we provide a comprehensive review of the existing literature in LLM-enhanced RL and summarize its characteristics compared to conventional RL methods, aiming to clarify the research scope and directions for future studies. Utilizing the classical agent-environment interaction paradigm, we propose a structured taxonomy to systematically categorize LLMs’ functionalities in RL, including four roles: information processor, reward designer, decision-maker, and generator. Additionally, for each role, we summarize the methodologies, analyze the specific RL challenges that are mitigated, and provide insights into future directions. Lastly, potential applications, prospective opportunities and challenges of the LLM-enhanced RL are discussed.
Index Terms:
Reinforcement learning (RL), large language models (LLM), vision-language models (VLM), multi-modal RL, LLM-enhanced RL.
I Introduction
Reinforcement learning (RL) is a powerful learning paradigm that focuses on control and decision-making, where an agent learns to optimize a specified target through trial and error interactions with the environment. The recent success of deep learning in classical domains, such as natural language processing (NLP) [1] and computer vision (CV) [2], has led to the rise of deep RL, which combines RL with high-capacity deep neural network approximators that enable agents to make decisions in complex environments. In the game realm, deep RL algorithms have achieved the beyond-human ability of decision-making in a multitude of arcade [3][4], real-time strategy [5][6], board [7][8], and imperfect information games [9][10]. Currently, the application of RL has to extend to real-world tasks in the fields of robotics [11], autonomous vehicles [12], healthcare [13], and conversational system [14].
However, deep RL algorithms face significant challenges when applied in real-world applications involving language and visual information, as the agent has to learn the features and a control policy jointly. To reduce the burden of visual feature learning, reference [15] decouples representation learning from RL. To handle language-involved tasks, a survey [16] calls for potential uses of NLP techniques in RL. Nevertheless, the capabilities of language models were limited at that time and the following four challenges still have not been addressed: 1) Sample inefficiency: deep RL agents require extensive interactions with the environment to learn effective policies, rendering them impractical for data collection scenarios that are expensive or risky [17][18][19]. 2) Reward function design: the performance of policy learning significantly depends on the design of the reward function [20]. Despite the fundamental importance, reward functions are known to be notoriously difficult to design [21][22] since they require a deep understanding of the task and often demand manual trial-and-error [23]. 3) Generalization: the generalization of deep RL agents remains daunting, as they frequently struggle to adapt to new, unseen environments, limiting the agents’ applicability in dynamic real-world settings [24][25]. 4) Natural language understanding: deep RL faces difficulties in natural language processing and understanding scenarios, where the nuances and complexities of human language present unique challenges that are not adequately addressed by current RL methodologies [26].
The recent emergence of large language models (LLMs) has marked a significant milestone in the field of NLP and shown various powerful capabilities in many real-world applications such as medicine [27], chemical [28] and embodied control in robots [29]. Compared to small language models, LLMs have emergent capabilities that are not present in small language models [30], such as in-context learning [31], reasoning ability [32] etc. Additionally, leveraging the vast amounts of training data, pre-trained LLMs are equipped with a broad spectrum of world knowledge [33]. Benefiting from these capabilities, the applications of language models have been shifted from language modeling to task-solving, ranging from basic text classification and sentiment analysis to complex high-level task planning [34] and decision-making [35][36][37].
With emergent capabilities, the potential of LLMs to address the inherent challenges of RL has recently gained popularity [38][39]. The capabilities of LLMs, particularly in natural language understanding, reasoning, and task planning, provide a unique approach to solving the above-mentioned RL issues. For sample inefficiency, reference [40] proposes a framework where LLMs can be employed to improve the sample efficiency of RL agents by providing rich, contextually informed predictions or suggestions, thereby reducing the need for extensive environment interactions. For reward function design, LLMs can aid in constructing more nuanced and effective reward functions, enhancing the learning process by offering a deeper understanding of complex scenarios [41]. For generalization, reference [42] proposes a framework that leverages language-based feedback for improving the generalization of RL policy in unseen environments. For natural language understanding, Pang et al.  [43] use LLMs to translate complex natural language-based instructions to simple task-specified languages for RL agents. These works show that LLM is a promising and powerful role that can contribute to the longstanding RL challenges.
Despite the advancements in the domain of integrating LLMs into the RL paradigm, there is currently a notable absence of comprehensive review in this rapidly evolving area. Additionally, though various methods are proposed to integrate LLMs into the RL paradigm, there is no unified framework for such integration. Our survey paper seeks to fill these gaps by providing an extensive review of the related literature, defining the scope of the novel paradigm called LLM-enhanced RL, and further proposing a taxonomy to categorize the functionalities of LLMs in the proposed paradigm.
I-A Contributions
This survey makes the following contributions:
  • LLM-enhanced RL paradigm: This paper presents the first comprehensive review in the emerging field of integrating LLM into the RL paradigm. To clarify the research scope and direction for future works, we define the term LLM-enhanced RL to encapsulate this class of methodologies, summarize the characteristics and provide a corresponding framework that clearly illustrates 1) how to integrate LLMs in classical agent-environment interaction and 2) the multifaceted enhancements that LLMs offer to the conventional RL paradigm.
  • Unified taxonomy: Further classifying the functionalities of LLMs in the LLM-enhanced RL paradigm, we propose a structured taxonomy to systematically categorize LLMs within the classical agent-environment paradigm, where LLMs are classified as information processors, reward designers, decision-makers, and generators. By such a categorization, a clear view of how LLMs integrate into the classical RL paradigm is offered.
  • Algorithmic review: For each role of LLM, we review emerging works in this direction and discuss different algorithmic characteristics from the perspective of capabilities. Based on this foundation, future applications, opportunities, and challenges of LLM-enhanced RL are analyzed to provide a potential roadmap for advancing this interdisciplinary field.
I-B Text Organization
The remaining sections are organized as follows. Section II provides foundational knowledge of both RL and LLM. Section III presents the concept of LLM-enhanced RL and provides its overall framework. Following this, Sections IVVVI, and 6 offer an in-depth analysis of LLMs within the RL context, exploring their roles as information processor, reward designer, decision-maker, and generator, respectively. Last, Section VIII discusses the application, opportunities and challenges of LLM-enhanced RL. Finally, Section IX concludes the survey.
II Background
In this section, we provide a concise overview of the classical RL paradigm and related challenges. Next, we explore the prevailing trend in RL—specifically, the fusion of multi-modal data sources, including language and visual information. Following this, we offer an introductory background on LLMs and outline the key capabilities that can enhance the RL learning paradigm.
II-A Background of Reinforcement Learning
II-A1 Classical Reinforcement Learning
In a classical RL paradigm shown in Fig. 1, the agent interacts with an environment through a trial-and-error process to maximize the specified rewards in the trajectory. In each step, the agent takes an action $a$ based on the observed state $s$ from the environment. By optimizing the policy $\pi$ (action controller), the agent maximizes the cumulative rewards. Such an optimization problem is usually formalized through the concept of Markov Decision Process (MDP), defined by the quintuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. Here, $\mathcal{S}$ denotes a set comprising all possible states, $\mathcal{A}$ denotes a set of all possible actions, $\mathcal{T}$ represents the state transition probability function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, $\mathcal{R}$ is a reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, and $\gamma$ (with $0\leq\gamma\leq 1$) is the discount factor. The objective in RL is to optimize the policy $\pi(a|s)$ such that the cumulative returns $\sum_{k=0}^{\infty}\gamma^{k}r_{k+1}$ is maximized.
Figure 1: Classical reinforcement learning paradigm.
II-A2 Challenges of Reinforcement Learning
While RL algorithms have made remarkable performance in recent years [3][7][44], there are still several longstanding challenges that limit the real-world applicability of RL:
  • Generalization in Unseen Environment: Generalization in unseen environments remains a significant challenge in the field of RL [45]. The core issue lies in the ability of RL algorithms to transfer learned knowledge or behaviors to new, previously unseen environments. RL models are often trained in simulated or specific settings, excelling in those scenarios but struggling to maintain performance when faced with novel or dynamic conditions. This limitation hinders the practical application of RL in real-world situations, where environments are rarely static or perfectly predictable. Achieving generalization requires models to not only learn specific task solutions but also to understand underlying principles that can be adapted to a range of situations.
  • Reward Function Design: Reward function is the principal contributing factor to the performance of an RL agent. Despite their fundamental importance, reward functions are known to be notoriously difficult to design, especially in contexts involving sparse reward environments and complex scenarios [22]. In sparse reward settings, where feedback is limited, reward shaping becomes essential to guide agents toward meaningful behaviors; however, this introduces the risk of inadvertently biasing the agent towards sub-optimal policies or overfitting to specific scenarios [46][47]. Conversely, for complex tasks, high-performance reward functions usually require massive manual trial-and-error since most designed rewards are sub-optimal [23] or lead to unintended behavior [48].
  • Compounding Error in Model-based Planning: Model-based RL is prone to the issue of compounding errors during planning. As the prediction horizon extends, errors in the model’s predictions accumulate, leading to significant deviations from optimal trajectories [49][50]. This problem is particularly acute in complex environments with high-dimensional state spaces. With their advanced predictive capabilities and understanding of sequential dependencies, LLMs could help mitigate these errors, leading to more accurate and reliable planning in model-based RL.
II-A3 Multi-modal Reinforcement Learning
With the advances in both CV and NLP, pattern recognition in vision and natural language has become increasingly powerful, and multi-modal data has been involved in the RL paradigm recently. Visual data is commonly involved in the observation space of RL when agents receive image-based information from the environment, e.g. in applications such as robots [51], video game control [3] etc. Compared to visual data, natural languages are usually included when RL agents are given specific tasks when interacting with the environments. The use of natural languages in RL can be divided into the following two categories [16]:
  • Language-conditional RL: In language-conditional RL, the problem itself requires the agent to interact with the environment through language. Specifically, there are two ways to integrate natural language in RL: 1) task description: the task or instruction is described in natural languages, e.g. instruction following, where the agents learn to interpret the instructions first and then execute actions; 2) action space or observation space: natural language is part of the state and action space, e.g.. text games, dialogue systems, and question answering (Q&A). This class of RL leverages natural language as a direct component of the RL process, guiding the agent’s actions and decisions within the language environment.
  • Language-assisted RL: In language-assisted RL, natural language is used to facilitate learning but not as a part of problem formulation. Two usages of language-assisted RL are: 1) communicating domain knowledge: the text containing task-related information can be helpful for agents. Therefore, wikis and manuals related to the environment can potentially assist the agents in such cases; 2) structuring policies: structuring policies is to communicate information about the state or dynamics of the environment based on language instead of representations of the environment or models. In such cases, language can be leveraged to shape representations towards a generalizable abstraction, such as using “avoid hitting the wall” instead of representations of a policy. This approach represents a more indirect use of natural language, serving as a guide or enhancer to the primary RL tasks.
The integration of multi-modal data challenges the RL paradigm since the agent has to simultaneously learn how to process complex multi-modal data and optimize the control policy in the environment [15]. Issues such as natural language understanding [52] and visual-based reward function design [53] require to be addressed.
II-B Background of Large Language Models
LLMs typically refer to the Transformer-based language models [54] containing hundreds of billions of parameters and being trained on massive text data (i.e., several terabytes (TB) scale) [55], such as GPT-3 [56], Galactica [57], PaLM [58] and LLaMA [59]. The vast number of parameters and internet-scale training data enable LLMs to learn a wide range of tasks, enabling more powerful abilities in language generation, knowledge representation, logical reasoning, etc., and better generalization capabilities on new tasks.
The development and effectiveness of LLMs are largely driven by Scaling Laws, i.e., as these models grow in size – both in terms of their parameter count and the data they are trained on – they tend to exhibit emergent abilities that are not present in small models [60][61][62], such as in-context learning, reasoning, and generalization. Here, we briefly introduce such capabilities of LLMs in detail:
  • In-context Learning: In-context learning capability eliminates the need for explicit model retraining or gradient update [56], as it can generate better responses or perform tasks by inputs cueing examples or related knowledge. Specifically, task-related texts are included in the prompts as context information, helping the LLMs to understand the situations and execute instructions.
  • Instruction Following: Leveraging diverse task-specific datasets formatted with natural language descriptions (also called instruction tuning), LLMs are shown to perform well on unseen tasks that are also described in the form of natural language [63][64]. Therefore, this capability equips LLMs with the ability to comprehend instructions for new tasks and effectively generalize across tasks not previously encountered, even in the absence of explicit examples.
  • Step-by-step Reasoning: For smaller models, tackling multi-step tasks, such as solving math word problems, often proves to be challenging. However, large language models can address the complex task effectively with sophisticated prompting strategies such as Chain of Thought (CoT) [65], Tree of Thought (ToT) [66], and Graph of Thought (GoT) [67]. These strategies structure the problem-solving process into sequential or hierarchical steps, facilitating a more articulated and understandable reasoning pathway. Additionally, prompts designed for planning enable LLMs to output sequences that reflect a progression of thoughts or actions, proving invaluable for tasks demanding logical sequence or decision-making outputs.
III Large Language Model-Enhanced Reinforcement Learning
Figure 2: Framework of LLM-enhanced RL in classical Agent-Environment interactions, where LLM plays different roles in enhancing RL.
III-A Definition
RL agents are often tasked with making robust and deliberate decisions using multi-modal information in real-real applications, whether in the MDP setting or within the context of specific task descriptions. Examples include robots designed to follow natural language instructions while navigating physical environments or visual games with tasks described in natural language [68][69][70]. However, it is challenging for conventional RL methods as the agent is required to simultaneously interpret complex multi-modal data and optimize control policies amidst ever-changing environments [71]. Compounding these challenges are issues like sample inefficiency, the difficulty of crafting reward functions that accurately reflect multi-modal inputs, and the need for robust generalization across varied tasks and settings.
The rapid advancements in LLMs present a viable solution to these challenges, thanks to their potent natural language understanding and reasoning abilities, coupled with recent progress in incorporating visual data processing [72]. This dual capability enables LLMs to interpret and act upon complex multi-modal information effectively, serving as a robust helper for enhancing the RL paradigm for real-world applications.
Nevertheless, despite the powerful functionalities of LLMs, the current studies are varied and lack a standard concept correctly specifying the systematic methodology, which impedes the advancement of research in this area. Therefore, we introduce the concept called LLM-enhanced RL as follows:
LLM-enhanced RL refers to the methods that utilize the multi-modal information processing, generating, reasoning, etc. capabilities of pre-trained, knowledge-inherent AI models to assist the RL paradigm.
Unlike traditional model-based reinforcement learning, a key characteristic of LLM-enhanced RL is that it utilizes the models with general knowledge, which means that the model has a considerable level of ability at the beginning of the learning process and better generalization compared to other data-driven models. In addition, the LLM-enhanced RL still uses the AI model and can continuously learn through the RL interaction data to improve its capabilities.
III-B Framework
The framework of LLM-enhanced RL is illustrated in the center of Fig. 2, which is founded on the classical agent-environment interaction paradigm. Along with the trial-and-error learning process, LLM processes the state information, redesigns the reward, assists in action selection, and interprets the policy after the action selection.
Specifically, on the one hand, when the agent receives the state and reward information from the environment, LLM is able to process or modify the information to either filter unnecessary natural language-based information or design appropriate rewards to accelerate the learning process, based on the natural language understanding and reasoning capabilities. On the other hand, when the agent is about to choose an action based on the observation, LLM can assist the action selection process by either simulating a world model or serving as the policy network to generate reasonable actions based on the modeling capability and common-sense knowledge. Additionally, after the action selection process, integrating state, reward, and action information, LLM can interpret the underlying possible reasons behind the policy selection, which helps human supervisors understand the scenarios for further system optimization.
Based on the functions of LLM in the framework, we extract characteristics of LLM-enhanced RL and further divide four different LLM roles in LLM-enhanced RL, including information processor, reward designer, generator, and decision-maker, which will be elaborated in the next subsections.
III-C Characteristics
The LLM-enhanced RL paradigm enhances the vanilla RL paradigm with the following characteristics:
  • Multi-modal Information Understanding: LLMs enhance RL agents’ comprehension of scenarios involving multi-modal information, enabling them to learn from tasks or environments described in natural language and vision data more effectively.
  • Multi-task Learning and Generalization: Benefiting from the multi-disciplinary pre-trained knowledge, LLMs empower RL agents by either designing rewards or generating reference actions, facilitating multi-task learning and generalization.
  • Improved Sample Efficiency: Given the inherent exploratory nature, the RL paradigm demands significant samples to learn. Pre-trained LLM can enhance data generation by simulation or leverage the prior knowledge to improve the sample efficiency of RL.
  • Long-Horizon Handling: RL becomes more challenging as the length of trajectory increases, due to the credit assignment problem. LLMs can decompose complex tasks down into sub-tasks to assist RL agents in planning over longer temporal horizons, aiding in the decision-making process for complex, multi-step tasks such as the Minecraft game.
  • Reward Signal Generation: Based on the context understanding and domain knowledge, LLMs contribute to the reward shaping and reward function designing, which help guide the RL towards effective policy learning in sparse-reward environments.
III-D Taxonomy
In this subsection, we illustrate the different roles of LLMs within the above framework, by detailing their functions and corresponding issues of RL they address:
  • Information Processor: When observation or task description involves language or visual features, it is challenging for the agent to comprehend the complex information and optimize the control policy simultaneously. To release the agent from the burden of understanding the multi-modal data, LLM can serve as an information processor for environment information or task instruction information by 1): extracting meaningful feature representations for speeding up network learning; 2) translating natural language-based environment information or task instruction information into formal specific task languages to reduce learning complexity.
  • Reward Designer: In complex task environments where the reward is sparse or a high-performance reward function is hard to define, using the prior world knowledge, reasoning abilities, and code generation ability, LLM can serve as two roles: 1) an implicit reward model to provide reward values based on the environment information, either by training or prompting; 2) an explicit reward model that generates executable codes of reward functions that transparently specifies the logical calculation process of reward scalars based on the environment specifications and language-based instructions or goals.
  • Decision-maker: RL faces the challenges of sample inefficiency and planning when exploring long-horizon environments. By serving as the decision-maker, pre-trained LLMs can perform: 1) direct decision-making: use the powerful sequence modeling ability and common-sense knowledge to improve the sample efficiency of offline RL; 2) indirect decision-making: act as an expert instructor that generates action candidates (high-level task planning) to shrink the action selection range or outputs reference policies to indirectly guide the updating directions of RL policies.
  • Generator: Model-based RL hinges on precise world models to learn accurate environment dynamics and simulate high-fidelity trajectories. Additionally, interpretability remains another important issue in RL. Using the multi-modal information understanding capability and prior common-sense reasoning ability, LLMs can be 1) a generator to generate accurate trajectories in model-based RL; 2) generate policy explanations with the prompts of related information in explainable RL.
IV LLM as Information Processor
The normal way for deep RL with language or visual information is to jointly process the information and learn a control policy, end-to-end. However, this demands the RL agent to learn to comprehend the information and manage the task simultaneously. Additionally, learning the language or visual features by simply relying on the reward function is challenging and may narrow the learned features to a narrow utility, hampering the generalization ability of the agent [15].
With the advances in unsupervised techniques and large-scale pre-trained models for CV and NLP, the decoupled structure, where the encoders are separately trained, has gained popularity [15][73][74]. Utilizing the powerful representation ability and prior knowledge, the pre-trained LLM or vision-language model (VLM) model can serve as an information processor to either extract an observation representation for downstream networks or translate the unbounded natural language to formal specification languages, which increase the sample efficiency and zero-shot performance of RL.
IV-A Feature Representation Extractor
Adopting the large pre-trained models in CV and NLP, the learned feature representation can be a scaffold embedding for downstream network learning and increase the sample efficiency. As illustrated in Fig. 3 (i), the usages can be further divided into two categories according to whether the model is trained simultaneously. One way is to directly use the frozen pre-trained model to extract embeddings from the observation $\mathcal{O}_{t}$ and another way is to further fine-tune the pre-trained model using contrastive learning with a contrastive loss $\mathcal{L}_{t}^{c}$ to achieve better adaptation in new environments.
Figure 3: LLM as an information processor.
IV-A1 Frozen Pre-trained Model
Using the frozen large-scale pre-trained model is the straightforward way. In reference [75], the author proposes History Compression via Language Models (HELM) to utilize a frozen pre-trained Language Transformer to extract history representation and compression and thus addresses the problem of partially observed MDP by approximating the underlying MDP with the past representation. Specifically, the framework first uses FrozenHopfield, a frozen associative memory to map observations $[o_{t-2},o_{t-1},o_{t}]$ to a compressed representation $h_{t}$ and then concatenate it with a learned encoding of the current observation via a convolutional neural network. Such a method solves the problem of how to effectively utilize the compressed history for policy optimization. After that, Semantic HELM [76] proposes a human-readable memory mechanism that summarizes past visual observations in human language and uses multimodal models to associate visual inputs with language tokens. The memory is a semantic database $\mathcal{S}$ constructed by encoding prompt-augmented tokens from the vocabularies of Contrastive Language-Image Pre-training (CLIP) [77] and the pre-trained language models. Given an observation $o_{t}$, the agent retrieves top-$k$ embeddings from $\mathcal{S}$ as actor-critic input to assist the policy optimization. Such a memory mechanism provides a human-readable representation of the past and helps the agent to cope with partially observable environments. However, one limitation of such frozen pre-trained models is that the representations cannot dynamically adjust according to the task and environment.
IV-A2 Fine-tuning Pre-trained Model
Some of the other literature modifies the pre-trained models with contrastive learning to learn an invariant feature representation, which improves the generalization capability.
Contrastive learning is a method that aims to learn representations from high-dimensional data by contrasting positive examples against negatives. Given a query $q$, the goal is to match query $q$ more closely to a positive key $k_{+}$ than to any negative keys $\mathbb{K}\setminus\{k_{+}\}$ in a set $\mathbb{K}$. This process is modeled using similarity measures, such as the dot product $(q^{T}k)$ or the bilinear product $(q^{T}Wk)$, where $W$ is a weight matrix. To effectively learn these representations, a contrastive loss function such as InfoNCE [78] is used:
$\mathcal{L}_{q}=\log\frac{\exp(q^{T}Wk_{+})}{\exp(q^{T}Wk_{+})+\sum_{k}^{K-1}\exp(q^{T}Wk_{i})}$ (1)
where $\exp$ is the exponential symbol and the whole $\mathcal{L}_{q}$ can be viewed as the log-loss of a softmax classifier, treating the matching of $q$ to $k_{+}$ as a multi-class classification problem among $K$ classes. By maximizing alignment between different changes of the same observation via the above loss, the model can learn the invariant representations, which addresses the adaptation problem induced by the variations of real-world environments and improves the sample efficiency in out-of-distribution generalization.
In reference [79], to achieve the zero-shot capability of embodied agents, the author devises a visual prompt-based contrastive learning framework that uses a pre-trained VLM to learn the visual state representations. By contrastively training the VLM on a pool of visual prompts along with the RL policy learning process, the learned representations are robust to the variations of environments and achieve competitive zero-shot capabilities. Based on the contrastive learning, another method ReCoRe [80] adds an intervention-invariant regularizer in the form of an auxiliary task such as depth prediction and image denoising to explicitly enforce invariance of learned representations to environment changes.
IV-B Language Translator
The unbounded and diverse representation of natural languages in both human instruction and environmental information impedes policy learning. LLM can be leveraged as a language translator to reduce the additional burden of comprehending natural language for RL agents and increase sample efficiency. As illustrated in Fig. 3 (ii), LLM transforms the diverse and informal natural language information into formal task-specific information, such as feature representation or task-specific languages, thus assisting the learning process of the RL agent.
IV-B1 Instruction Information Translation
One application of LLM is to translate natural language-based instructions for instruction-following applications. In reference [43], the author investigates an inside-out scheme for natural language-conditioned RL by training an LLM that translates the natural language to a task-related unique language. Such an inside-out scheme prevents the policy from being directly exposed to natural language instructions and helps efficient policy learning. Another literature STARLING [81] uses LLM to translate natural language-based instructions to game information and example game metadata. The translated information is then fed forward to Inform7, an interactive fiction game engine, to develop a large amount of text-based games for RL agents to master the desired skills.
IV-B2 Environment Information Translation
On the other hand, LLM can also be used to translate the natural language environment information into formal domain-specific language that can specify MDP information, which converts natural language sentences into grounded usable knowledge for the agent. Previous works generally ground natural language into individual task components such as task objectives description [82], rewards [83] and polices [84][85]. To unify the information about all the components of a task, [86] introduces RLang, a grounded formal language capable of expressing information about every element of an MDP and solutions such as policies, plans, reward functions and transition functions. By using LLM to translate natural language to RLang and train RL agents upon RLang, the agents are capable of leveraging information and avoid having to learn tabula rasa.
IV-C Summarization and Outlook
For information processing, LLM is used to accelerate RL learning processing by decoupling the information processing task and the controlling task, where LLM extracts feature representations or handles the natural language-based information. In the following, we list potential directions for future research.
  • Feature Representation Extractor: Although the use of LLMs as feature representation extractors has shown promise in enhancing RL, several challenges persist in future research. First, the computational overhead of LLMs is large for RL agents, an efficient and simplified feature extractor without excessive computational demand is required. Second, although LLMs can be used to learn features from high-dimensional inputs, their representation may lack generalization and be vulnerable to the high diversity inherent in environments or not take into account specific characteristics for solving control tasks. Future work may consider exploiting the task compositionality to increase the generalization of learned representations.
  • Language Translator: Current researches only focus on language translation and directly trains LLM to translate the natural language to task-related language while not considering the more generalization problem where the environment or task information is multi-modal. Future work may consider adding the multi-modal translation ability to achieve a more general translator and help agent learning.
V LLM as Reward Designer
The reward signal is the most important information to instruct agent learning in RL [22]. However, despite the fundamental importance, high-performing reward functions are known to be notoriously difficult to design [87]. On the one hand, specifying human notions of desired behavior is difficult via designed reward functions or requires huge expert demonstrations. On the other hand, dense rewards that accurately provide learning signals require either manually decomposing the general goal into sub-goals [88] or rewarding interesting auxiliary objectives [89]. Nevertheless, both of these methods suffer from the need for expert input and meticulous manual crafting [23].
Benefiting from pre-trained common-sense knowledge, code generation, and in-context learning ability, LLM has the potential to design or shape reward functions for DRL by leveraging natural language-based instructions and environment information. In this section, we review recent literature in which LLMs act as reward models that implicitly provide reward values or explicitly write executable reward function codes detailing the calculation process of reward scalars.
Figure 4: LLM as a reward designer.
V-A Implicit Reward Model
A large pre-trained model can be an implicit reward model that directly provides auxiliary or overall reward value based on the understanding of task objectives and observations. The methods are illustrated in Fig. 4 (i). One way is by directly prompting with language descriptions and another way is by scoring the alignment between the feature representation of the visual observations and language-based instructions.
V-A1 Direct Prompting
Reference [90] simplifies the manual reward design by prompting an LLM as a proxy reward function with examples of desirable behaviors and the preferences description of the desired behaviors. Reference [91] proposes a Read and Reward framework that utilizes LLMs to read instruction manuals to boost the learning policies of specific tasks. The framework includes a QA extraction module for information retrieval and summarization and a Reasoning module for evaluation. Experimentally, they show RL algorithms by their design can obtain significant improvement in performance and training speed. In reference [39], Carta et al. proposes an automated reward shaping method where the agent extracts auxiliary objectives from the general language goal. Using a question generation (QG) and question answering system (QA), the framework guides the agent in reconstructing partial information about the global goal and provides an intrinsic reward signal for the agent. This intrinsic reward incentivizes the agent to produce trajectories that help them reconstruct the partial information about the general language goal. To acquire a generalizable policy by continually learning a set of tasks is challenging for RL agents since the agent is required to retain the previous knowledge while quickly adapting to new tasks. Reference [92] introduces the Lafite-RL (Language agent feedback interactive Reinforcement Learning) framework that provides interactive rewards mimicking human feedback based on LLMs’ real-time understanding of the agent’s behavior. By designing two prompts, one to let LLM understand the scenario and the other to instruct it about the evaluation criterion, their framework can accelerate the RL process while freeing up human effort during the interaction between agent and environment.
V-A2 Alignment Scoring
For visual RL, some literature utilizes vision-language models as reward models to align multi-modal data and calculate the similarity score using metrics such as cosine similarity. Rocamonde et al. employs the CLIP model as a zero-shot reward model to specify tasks via natural language [53]. They first compute the probability $p_{o_{t},l}$ that the agent achieves a goal given by language description $l$ out of a set of potential goals $l^{\prime}\in\mathcal{L}$ in the task set $\mathcal{L}$ using the softmax computation with temperature $\tau$ over the cosine similarity between visual state embeddings $f_{\theta}(o_{t})$ and language description embeddings $g_{\theta}(l)$ across the set of potential goals $l^{\prime}$:
$p_{o_{t},l}=\frac{\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l)/\tau)}{\sum_{l^{\prime}}\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l^{\prime})/\tau)}$ (2)
Then the reward is obtained by a binary reward function $r_{t}=r(o_{t+1},l)=\mathbb{I}[p_{o_{t+1},l}\geq\beta]$, which thresholding the probability. Their framework only requires a single-sentence text prompt description of the desired task with minimal prompt engineering. Another work [93] constructs reward signals based on the similarity between natural language-based description and embeddings from the pre-trained VLM encoder. By labeling the expert demonstration with the reward signals, the framework effectively mitigates the problem of mis-generalization. In reference [94], the authors propose a Language Reward Modulated Pretraining (LAMP) framework by leveraging the zero-shot capabilities of vision-language models to generate noisy, albeit shaped exploration rewards. Such rewards are generated by computing the contrastive alignment between language instructions and the image observations of the agent in the pre-training environment. Using the LAMP framework, the generated rewards are optimized with standard novelty-seeking exploration rewards for language-conditioned policy and can warm start sample-efficient learning on robot manipulation tasks.
V-B Explicit Reward Model
Another way to design reward functions is by generating executable codes that explicitly specify the details of the calculation process, as illustrated in Fig. 4 (ii). Compared to the implicit reward value provision, this explicit way transparently reflects the reasoning and logical process of LLMs and thus is readable for humans to further evaluate and optimize.
To help robots learn low-level actions, reference [95] harnesses the code generation ability of LLMs to define the lower-level reward parameters based on high-level instructions. Using such a reward design paradigm, their work bridges the gap between high-level language instructions to low-level robot actions and can reliably tackle 90% of the designed tasks compared to the 50% of the baseline. Motivated by the capability of LLM for self-refinement [96], reference [97] proposes a framework with a self-refinement mechanism for automated reward function design, including initial design, evaluation and self-refinement loop. Their results indicate that the LLM-designed reward functions are able to rival or surpass manually designed reward functions. Similarly, Eureka [98] develops a reward optimization algorithm with self-reflection. The algorithm uses an environment source code and task description to sample reward function candidates from a coding LLM, evaluate them over RL simulation, and use the result to progressively improve the reward outputs. The results show that their proposed method can achieve human-level performance on reward design and solve dexterous manipulation tasks that were previously infeasible by manual reward engineering. Another work Text2Reward [99] generates shaped dense reward functions as executable programs based on the environment description. Given the sensitivity of RL training and the ambiguity of language, the RL policy may fail to achieve the goal. Text2Reward addresses the problem by executing the learned policy in the environment, requesting human feedback and refining the reward accordingly.
V-C Summarization and Outlook
The use of LLMs as reward designers in reinforcement learning offers a more natural and efficient way to design complex reward functions. By leveraging natural language processing capabilities, LLMs simplify the traditionally challenging task of reward function design, enhancing both the efficiency and effectiveness of RL algorithms. In the future, we expect advancements in the following areas:
  • Implicit Reward Model: Future work may consider focusing on improving the alignment of LLM-generated rewards with complex human intentions and behaviors. This involves refining the quality of language instructions to reduce ambiguities and inaccuracies, ensuring that the reward functions align precisely with human notions of desired behavior. Additionally, generalization and transferability of LLM-generated rewards across different tasks and environments, especially in complex, high-dimensional visual environments, is also an important direction to explore.
  • Explicit Reward Model: A key limitation in reward code generation is its dependency on pre-trained common-sense knowledge, which can be restrictive for highly specialized tasks not covered in training data. Enhancing prompts with detailed, task-specific information is essential for effectiveness. Additionally, the reliance on manually designed templates for motion descriptions limits the adaptability, suggesting a need for future research focused on automating or unifying template design processes.
VI LLM as Decision-Maker
LLMs trained on a massive amount of data show impressive results in language understanding tasks [56], instruction-following [100], vision-language navigation [101] and tasks requiring planning and sequential reasoning [102]. Such success motivates researchers to explore the potential of LLMs for decision-making problems as a direct decision-maker that generates actions or an indirect decision-maker that instructs the actions. In direct decision-making, Transformer-based models such as Decision Transformer [103] show great potential in offline RL when dealing with sparse-reward and long-horizon tasks [104]. As LLM itself is a large-scale Transformer-based model, LLM can be viewed as an enhanced Decision Transformer with more pre-trained world knowledge and powerful modeling abilities, which unleash the power of pre-trained LLMs for offline RL [105]. In indirect decision-making, LLMs can be prompted to generate “expert” actions that instruct the action selection of RL agents.
Figure 5: LLM as a decision-maker.
VI-A Direct Decision-Maker
In direct decision-making, LLM enhances the Decision Transformer-based methods with a more powerful pre-trained model with more prior world knowledge to solve sparse-reward and long-horizon tasks with a higher sample efficiency.
As shown in Fig. 5 (i), Decision Transformer-based methods directly predict the actions from the perspective of sequence modeling by framing the learning process of offline RL into a supervised learning problem. Given each trajectory $\tau$ as a sequence of ordered return-to-go $\hat{R}$, action $a$, and states $s$ as follows:
$\tau=($ $\hat{R}_{t_{0}},s_{t_{0}},a_{t_{0}},\hat{R}_{t_{0}+1},s_{t_{0}+1},a_{t_{0}+1},\ldots,$
$\hat{R}_{t_{0}+K-1},s_{t_{0}+K-1},a_{t_{0}+K-1})$ (3)
where $K$ is the context length, the model learns to predict the future action $a^{\prime}_{t}$ by minimizing the squared error loss:
$\mathcal{L}_{decision}=\sum_{t=t_{0}}^{t_{0}+K-1}\left\|a_{t}-a_{t}^{\prime}\right\|_{2}^{2}$ (4)
Reference [106] investigates the transferability of pre-trained sequence models on vision and language domains when fine-tuned on other offline tasks such as control and games. Results show that using Wikipedia-pre-trained and GPT2 language models, there are consistent performance gains in terms of reward and convergence speed in various environments. Reference [107] proposes to use pre-trained LLM as a general scaffold for task-specific model learning across various environments by adding goals along with observations as the input and converting them to sequential data. Experiments demonstrate that language modeling improves combinatorial generalization in policy learning and can substantially improve out-of-distribution performance in new tasks. To unify language reasoning with actions in a single policy, reference [108] generates textual captions interleaved with actions when training the Transformer-based policy. Results show that by using captions describing the next subgoals, the reasoning policy can consistently outperform the caption-free baseline. In-context RL is one approach to generalist RL by recalling the memory from past experiences. Reference [109] redesigns the offline in-context approach to successfully train long-sequence Transformers over entire rollouts to tackle the challenges of generalization, long-term memory, and meta-learning. To deal with situations when data is scarce and the environment is risky, reference [105] proposes Language Models for Motion Controls (LaMo) to employ pre-trained LLMs with LoRA fine-tuning method that augments the pre-trained knowledge with in-domain knowledge. The experiments indicate LaMo achieves state-of-the-art performance in sparse-reward tasks and has performance gains in data-limited scenarios. To integrate multi-modal data, e.g., vision and language, into the offline RL, reference [110] co-fine-tunes vision-language models on both robotic trajectory data and Internet-scale vision-language tasks, e.g., visual question answering. In the framework, they incorporate the actions as natural language tokens and co-fine-tune models on both vision and language datasets. Results show that such co-fine-tune methods can increase generalization performance and the chain of thought reasoning can help the agent perform multi-stage semantic reasoning and solve complex tasks.
VI-B Indirect Decision-Maker
The indirect decision-maker role is illustrated in Fig. 5 (ii). As an indirect decision-maker, LLM instructs the action selection by either generating a set of action candidates or providing a reference policy that instructs or regularizes policy learning. By instructing the action choices, LLM addresses the challenges posed by enormous action spaces and natural language in applications such as instruction-following and text-based games.
VI-B1 Action Candidates
When LLM generates a set of action candidates, RL agents can further re-rank candidates based on the value function and maximize the cumulative rewards. In text-based games, action spaces are large and only a tiny fraction of action commands are accessible in any given game state. To tackle such problems, reference [111] proposes the Contextual Action Language Model (CALM) framework to generate a set of action candidates for each state in a text-based game. By training the language models on human gameplay based on the game history, the language model generates action candidates for another network to calculate the Q-values and select the optimal action. Following this work, SayCan [102] extends the application into the instruction-following agent in the robotic field. In their work, the agent integrates an LLM to generate high-level action planning and then execute low-level skills. Different from CALM, the LLM is not trained and the action candidates are generated based on prompting with the history and current observation.
VI-B2 Reference Policy
When LLM provides a reference policy using LLM, the policy updating process is modified based on the referenced policy. In human-AI collaboration, RL often converges to the equilibrium that humans may not prefer, especially in settings where human quality human data is rarely available. To solve such problems, instructRL [112] uses LLM to generate prior policy based on human instructions and uses this prior to regularizing the RL objective. Specifically, instructRL augments the policy update function with an auxiliary term $p_{\mathrm{LLM}}[\mathrm{lang(}a_{t}\mathrm{)|lang(}\tau_{t}^{i}\mathrm{),inst}]$, the probability of choosing an action based on the trajectory and instructions. Experiments show instructRL converges to policies aligned with human preferences. To solve the adaptation problem of LLM-based agents in real-time dynamic environments, Zhou et al. [113] trains a smaller-scale specialized student agent using the instruction from an LLM-based teacher agent and gradually distills the prior knowledge to the local student model, resulting in less data for training a student agent.
VI-C Summarization and Outlook
As a decision-maker, LLM generates the action directly or indirectly based on observation and instructions. In direct decision-making, the sequence modeling capability of LLM is employed to improve the sample efficiency and generalization benefiting from the pre-trained world knowledge in LLM. For indirect decision-making, LLM guides the decision-making by generating action candidates or instructing the policy updating directions. In the following, challenges and future directions of the two roles are listed below:
  • Direct Decision-maker: Directly employing pre-trained large-scale LLM as the policy network demands huge computational resources and huge data for fine-tuning it in specific tasks. Future work may consider more cost-effective methods such as the Low-Rank Adaptation (LoRA) [114] to exploit the power of LLM in direct decision-making.
  • Indirect Decision-maker: Since the LLM acts as an instructor to provide reference policy, the bias and limitations are also inherited by the agent. In addition, LLM itself cannot inherently interrogate or intervene in the environment, which limits LLMs’ capabilities to intentionally aggregate information. Therefore, how to use the information gained from real-world interactions to improve the LLM itself in terms of actuality and reasoning is another important problem.
VII LLM as Generator
The generative capability of LLMs can be applied to environmental simulation and behavior explanation. On the one hand, growing interests in applying RL to the real world and risky data-collection process suggests a need for imaginary rollouts [115][116]. Possessed with a powerful modeling capability and world knowledge, LLMs can serve as a world model simulator to learn complex environmental dynamics with high fidelity by iteratively predicting the next state and reward, thus increasing the sample efficiency in model-based RL [117][118]. On the other hand, in RL, interpretability remains an important security issue in current black-box AI systems as they are increasingly being deployed to help end-users with everyday tasks. Explanations of policy can improve the end-user understanding of the agent’s decision-making and inform the reward function design for agent learning. In such aspects, LLM can act as a policy interpreter based on their knowledge and reasoning ability. In this section, we classify the above two roles of LLM as a generator and review recent related works.
Figure 6: LLM as a generator.
VII-A World Model Simulator
Serving as a world model simulator, LLM is trained as a 1) trajectory rolloutor, which auto-regressively generates accurate trajectories for the agent to learn and plan; 2) dynamics representation learner, which predicts the latent representation of the world using representation learning. A flow chart of the world model is illustrated in Fig. 6 (i). From the real world, knowledge and real data can be collected to construct a world model simulator, which further models the dynamics representation of the world, generates trajectories and helps the policy learning of the agent in the real world.
VII-A1 Trajectory Rolloutor
Similar to the decision transformer [119] in offline RL, pre-trained large-scale models are used in model-based to synthesize trajectories. In 2022, Micheli et al. proposed IRIS, an agent that employs a discrete autoencoder and an autoregressive Transformer to learn the world model for Atari games [120]. With the equivalent of two hours of gameplay in the Atari 100k benchmark, the proposed method outperforms humans in 10 out of 26 games. Similarly, reference [121] applies a transformer to build a sample-efficient world model for Atari games. Utilizing such a Transformer-based world model (TWM), the RL agent can solve the long-term dependency and train a state-of-the-art policy on the Atari 100k benchmark based on generated meaningful experiences from TWM. Visual RL enables the RL agent to learn from visual observations effectively. Reference [122] propose TransDreamer, a Transformer-based model-based RL agent that leverages a Transformer for dynamics predictions in 2D and 3D visual RL tasks. Experiments show the TransDreamer agent can outperform Dreamer with long-range memory access and memory-based reasoning. Based on the development of supervised pre-training methods, Seo et al. proposes a framework that learns world model dynamics from action-free video representations [123]. The framework adds a video-based intrinsic bonus for better-guiding exploration to effectively encourage agents to learn diverse behaviors. The experimental results demonstrate their proposed method can improve the performances of vision-based RL on manipulation and locomotion tasks by transferring the pre-trained representations from unseen domains.
VII-A2 Dynamics Representation Learner
Using representation learning techniques, the latent representation of the future can be learned to assist decision-making. Reference [124] introduces a visual model-based RL framework that decouples visual representation learning and dynamic learning by training an autoencoder with a vision Transformer to reconstruct pixel-given masked observations and learn the dynamics from the latent space. By such a decoupling approach, their proposed method achieves state-of-the-art performance on visual robotic tasks from Meta-world and RLBench. Utilizing the fact that language contains rich information signals and can help agents to predict the future, Lin et al. proposes Dynalang [40], where an agent that learns a multi-modal world model to predict future text and image representations and learns to act from imagined world representation. Different from other works that use language only for predicting actions, Dynalang can acquire rich language understanding by using past language and also to predict future language, videos and rewards. To solve the out-of-distribution generalization problem in visual control tasks of RL, reference [125] proposes the Language Grounded World Model (LanGWM), which focuses on learning language-grounded visual features to enhance the world model learning. To improve the generalization of the learned visual features, they masked the bounding boxes and predicted them with given language descriptions. Utilizing the expressing ability of language in higher-level concepts and global contexts, the proposed LanGWM method yields state-of-the-art results on out-of-distribution tests.
VII-B Policy Interpreter
Explainable RL (XRL) is an emerging subfield of both explainable machine learning and RL that has attracted considerable attention recently. XRL aims to elucidate the decision-making process of learning agents. According to a survey in explainable RL [126], the categories of XRL include the feature importance, learning process and MDP, and policy level. Currently, the usage of LLMs in XRL has only been limited to the policy level, i.e., as the policy interpreter. Therefore, though there is limited literature related to this area, we think this field is important and requires more attention in the future. The following will introduce the role of LLMs as policy interpreters and an outlook regarding other categories of XRL will be provided in the last subsection.
As the policy interpreter, LLM generates explanations with the prompts of state and action description or trajectory information. An illustration is depicted in Fig. 6. Using the trajectory history of states and actions as context information, LLMs can be prompts to generate readable interpretations of current policies or situations for humans.
Das et al. [127] proposes a unified framework called State2Explanation (S2E), that learns a joint embedding model between state-action pairs and concept-based explanation. Based on the learned models, the explanation can help inform reward shaping during an agent’s training and provide insights to end-users at deployment. Another work  [128] first distills the policy into a decision tree, derives the decision path, and then prompts an LLM to generate a natural language explanation based on the decision path.
VII-C Summarization and Outlook
As a generator, LLMs can be integrated into model-based RL or explainable RL, i.e., serving as world model simulators or policy interpreters, respectively. As world model simulators, LLMs enhance model-based RL by auto-regressively generating accurate trajectories (trajectory rollout) and by predicting latent world representations (world representation learners), significantly improving sample efficiency and decision-making accuracy. In the realm of explainable RL, LLMs provide valuable insights for both end-user understanding and reward shaping by generating explanations based on trajectory information. However, the current usages of LLMs in explainable RL are rather limited and have great potential for future work. Below are discussions on the key limitations and future work directions of the two roles:
  • World Model Simulator: LLMs encounter challenges in aligning their abstract knowledge with the specific requirements of different environments, leading to limitations in functional competence and grounding. This misalignment affects their effectiveness in generating trajectories and interacting with the environment. Additionally, the current model-based agents, which rely on purely observational world models, present difficulties for human adaptation. These models are typically modifiable only through observational data, which is not an effective means for humans to communicate complex intentions or adjustments. Future work may consider integrating language instructions or an adapter into the world model to achieve a flexible and adaptive world model.
  • Policy Interpreter: As policy interpreters, the quality of explanations depends on the LLM’s understanding of feature representations and implicit logic of policy. How to utilize domain knowledge or examples to improve the understanding of a complex policy is still a major issue. From the field of explainable RL, LLM may unlock its potential in different aspects. For example, from the view of feature importance, LLMs can be leveraged to interpret the correlation between observations and policy selection; from the view of the learning process and MDP, the agent’s behavior can be influenced by the MDP (such as reward $R$ and episode length $T$) or interactions with the MDP, where such influences can be revealed by asking the LLM to answer “why” and “why-not” questions with MDP context information.
VIII Discussion
In previous sections, the concept ”LM-enhanced RL” and a corresponding framework have been developed and extended to the integration of multi-modal AI models such as visual-language models. Further, the different roles of LLMs in enhancing the RL paradigm have been discussed. In this section, we analyze potential applications of the LLM-enhanced RL paradigm and then discuss its future opportunities and challenges considering the untapped capabilities and limitations of LLMs, particularly in environments with multi-modal information.
VIII-A Applications of LLM-Enhanced RL
Based on the characteristics of LLM-enhanced RL, such as multi-modal information understanding and multi-task learning and generation, we believe that LLM-enhanced RL opens up a wide array of potential applications. Here we list several applications to inspire researchers.
  • Robotics: RL is widely used in robots to learn how to make decisions and execute actions to achieve goals. Utilizing natural language understanding and general logical reasoning abilities, LLM-enhanced RL can 1) improve the efficiency of human-robot interaction, 2) help robots understand human needs and behavioral logic, and 3) enhance decision-making and planning capabilities.
  • Autonomous Driving: Autonomous driving uses RL to make decisions in complex, ever-changing environments that involve understanding both sensor data (visual, lidar, radar) and contextual information (traffic laws, human behavior). LLM-enhanced RL could employ LLMs to 1) process this multi-modal information and natural language instructions; or 2) design comprehensive rewards based on multi-disciplinary metrics such as safety, efficiency, and passenger comfort.
  • Energy Management: In the energy system, operators or users apply RL to efficiently manage the usage, transportation, conversion and storage of multiple energy with high uncertainty brought by renewable resources. LLM-enhanced RL in such cases can 1) improve the RL agents’ ability to handle multi-objective tasks, such as economy, safety, and low carbon, by reward function designing, and 2) increase the sample efficiency for the new energy system.
  • Healthcare Recommendation: RL is used to learn recommendations or suggestions in healthcare [13]. LLMs can utilize the domain knowledge to analyze the vast amount of patient data and medical histories, therefore, 1) accelerating the learning process of RL recommendation and 2) providing more accurate diagnostic and treatment recommendations in healthcare.
VIII-B Opportunities for LLM-Enhanced RL
Although current works in LLM-enhanced RL have already shown better performance in several aspects, more unexplored areas remain to be explored and may lead to significant improvement. Below, we summarize the potential opportunities of LLM-enhanced RL from both the perspectives of RL and LLM capabilities, respectively.
  • RL: Existing work such as [112][108][107] mainly focus on the general RL while various specialized branches of RL are still under-exploited, e.g. multi-agent RL, safe RL, transfer RL, explainable RL, and other niches within the field. In multi-agent RL, LLM-enhanced strategies could be employed to facilitate communication and collaboration among RL agents. The natural language understanding capabilities of LLMs can be used to interpret and generate instructions or strategies among agents, enhancing cooperative behaviors or competition strategies; Safe RL could benefit from the reasoning and predictive capabilities of LLMs to design cost functions that encourage safety criteria compliance, reducing the risks of dangerous exploration; In transfer RL, LLMs can assist in identifying similarities between tasks or environments, leveraging their vast knowledge base to facilitate knowledge transfer and thus improve learning efficiency and adaptability across different tasks.
  • LLM: Present methodologies such as [102][39][90] predominantly harness LLMs by employing straightforward prompting techniques to elicit responses. However, integration of LLMs with the tool manipulation capability unlocks new possibilities of LLMs. Tools such as mathematical solver and internet browser, retrieval-augmented generation (RAG) techniques, and API assistance can be leveraged to equip LLMs with a plethora of external knowledge and computational capabilities, improving the performance of LLM-enhanced RL in tasks that require real-time data fetching, external knowledge and mathematical model solving.
VIII-C Challenges of LLM-Enhanced RL
While LLM improves different issues in the RL paradigm, the success of LLM-enhanced RL is inherently tied to the capabilities and limitations of the underlying LLM. The primary concerns revolve around the dependency on LLM capabilities, interaction efficiency, and the broader implications of deploying such systems in real-world scenarios.
  • Dependence on LLM’s Capabilities: The LLM-enhanced RL paradigm depends on the capabilities of LLMs, which affects the performance of RL agent. Biases and hallucination inherent in pre-trained LLMs can lead to skewed or inaccurate interpretations of multi-modal inputs. Additionally, despite the vast knowledge base of LLMs, they may struggle to adapt to highly specific or novel task environments not well-represented in their training data. Therefore, to effectively integrate LLMs into RL paradigm, a comprehensive evaluation is required in advance.
  • Interaction Efficiency: The integration of LLMs into the RL learning process introduces complexities in terms of computational overhead and response times, which slow down the learning process of RL. These complexities require advances in both algorithms and hardware implementation for faster inference with fewer computational resources.
  • Ethical, Legal, and Safety Concerns: In practical usages, the use of LLMs involves complex ethical, legal, and safety concerns. Data privacy, intellectual property, and accountability for AI decisions should be carefully discussed. In addition, LLMs in RL introduce new sectors for adversarial attacks, potentially compromising the integrity and safety of the system.
IX Conclusion
LLMs, with their pre-trained knowledge bases and powerful capabilities such as reasoning and in-context learning, present as a viable solution to enhance RL in terms of natural language understanding, multi-task generalization, task planning, and sample efficiency. In this survey, we have defined this paradigm as LLM-enhanced RL and summarized its characteristics, together with opportunities and challenges. To formalize the research scope and methodology of LLM-enhanced RL, we propose a structured framework to systematically categorize the roles of LLM based on the functionalities within the classical agent-environment interaction paradigm. According to functionalities, we categorize the roles of LLMs into information processor, reward designer, decision-maker, and generator. For each category, we review current literature based on their methods and applications, outline the methodologies, discuss the addressed issues, and provide insights into future directions. These are listed below:
  • Information Processor: LLMs extract observational representations and formal specification languages for RL agents, thus increasing the sample efficiency. Future directions include incorporating goal-based information and integrating multi-modal environmental information to obtain stronger information extraction ability.
  • Reward Designer: For intricate or un-quantifiable tasks, LLMs can leverage pre-trained knowledge to generate high-performing rewards that are notoriously difficult for humans to design. Current methods still require consistently modifying the prompts and instructions of LLMs, calling for future work on automated self-evolving frameworks without human intervention.
  • Decision-Maker: LLMs generate direct actions or indirect advice for the agent to improve exploration efficiency. Huge computational overhead is a major issue in online interaction. Cost-effective methods to reduce the huge computational overhead of LLM in online RL is an important direction.
  • Generator: With the generative capability and world knowledge, LLMs are used as 1) a high-fidelity world model to reduce real-world learning cost; and 2) a language-based policy interpreter to explain the agent policy. Leveraging human instructions to improve the accuracy and generalizability of the world model and policy interpreter would be a crucial direction.